# Causal Mechanism Estimation in Multi-Sensor Systems Across Multiple Domains

### v3

@@ II-B Continuous Optimization-based Methods
$\begin{gathered}\min_{\theta}\quad\frac{1}{n}\sum_{j=1}^{d}\ell(\mathbf{x}_{j},\mathrm{MLP}(\mathbf{X};\theta_{j}))+\lambda\|A_{j}^{(1)}\|_{1,1}\\ \text{subject to }h(W(\theta)):=\mathrm{tr}(e^{W(\theta)\circ W(\theta)})-d=0~,\end{gathered}$ (7)
where $\ell$ is a loss function such as the least squares or the negative log-likelihood, $\ell_{1}$ regularization $\|A_{j}^{(1)}\|_{1,1}$ is used to enforce sparsity in the learned DAG, and the acyclicity constraint is enforced by $h(W(\theta))$ where $\circ$ denotes the Hadamard product and $e^{A}$ is the matrix exponential of AA. Using the machinery of augmented Lagrangian method [24], the solution of the equality-constrained problem in (7) can be well approximated by solving an unconstrained $\ell_{1}$-penalized smooth minimization problem with the L-BFGS-B algorithm [25].
@@ III Methodology
III-A Step 11: Causal Discovery on Pooled Data
@@ 4   Update $\theta$ to minimize the loss
III-B Step 22: Detection of Stable Variables
@@ 8      Add node $X_{j}$ to $X_{\mathrm{stable}}$
III-C Step 33: Causal Discovery on Individual Datasets
@@ IV Experiments
In this section, we generate synthetic data comprised of multiple domains based on linear Gaussian models to compare the performance of CICME and baseline methods. To assess CICME’s ability to detect common causal mechanisms, we consider 44 scenarios inspired from a real manufacturing process where causal mechanisms may shift across domains.
@@ IV-A Baselines
  • Applying NOTEARS-MLP on the pooled dataset, referred to as NOTEARS-pool. It is equivalent to step 11 of CICME.
  • Applying NOTEARS-MLP in each domain individually, referred to as NOTEARS-ind. It is equivalent to step 33 of CICME with $X_{\mathrm{stable}}$ being an empty set.
@@ IV-E Results
We first examine the capability of CICME to identify common causal mechanisms across all domains based on the HSIC test results from step 22. Note that in (E1) and (E2), $X_{4}$ exhibits changing causal mechanisms due to varying values of $H$ across domains, while in (E3) and (E4), $X_{2}$ is the expected unstable variable due to different distributions of $N_{2}$ across domains. Table I summarizes the total number of cases (out of $100$) in which each variable is identified as a stable variable. Additionally, the LSHD is averaged over those cases to indicate the accuracy of the estimated local structure. It can be seen that, when provided with $1000$ samples, CICME is able to reliably identify variables with changing causal mechanisms, as evidenced by the rare cases of misclassifications highlighted in bold. However, they become more frequent with a reduced sample size, since fewer samples provide less information about the population, leading to less statistical power of the HSIC test. Similarly, the accuracy of the estimated causal parents of stable variables also deteriorates when provided with fewer samples, as illustrated by the higher LSHD. This may adversely impact the performance of CICME by carrying over inaccurate common causal mechanisms to step 33.
In addition, we also observe that stable variables are occasionally identified as unstable. While this leads to higher degrees of freedom in updating the weights of the model, and hence, more training effort in step 33, it typically does not affect the overall performance. However, an exception is observed in (E4), where $X_{3}$ is identified as unstable significantly more often than stable even for $1000$ samples, despite the ground truth causal mechanism of $X_{3}$ being the same across all domains (i.e., $X_{3}=X_{1}+X_{2}+N_{3}$). We suspect that this is due to convergence issues in the constrained loss function. This can be hypothetically solved by finetuning the MLP based on the least squares loss before applying the HSIC test in step 22.
@@ IV-E1 Causal Structure
Moreover, Fig. 3 compares the capability of all methods to recover causal structures when the causal mechanism of $X_{2}$ changes across domains. When $X_{2}$ follows a distribution with shifted means across domains, the performance of CICME is on par with either of the baseline methods, as shown in Fig. 3(a). Similarly, NOTEARS-pool again outperforms NOTEARS-ind for $10$ samples, for the same reason mentioned earlier. However, this trend changes as soon as $X_{2}$ is fixed in each domain. Fig. 3(b) illustrates that NOTEARS-pool significantly outperforms NOTEARS-ind regardless of the sample size. This expected behavior can be attributed to the variance of $X_{2}$ in the observed data: The variance of $X_{2}$ equals to zero in each individual dataset and becomes non-zero in the pooled dataset, making it possible to infer the causal relationship between $X_{2}$ and other variables. Due to the pooling operation in step 11, CICME is also able to outperform NOTEARS-ind regardless of the sample size, and it even achieves the best performance among all when the sample size is $10$. However, for a sample size of $1000$, CICME fails to match the performance of the best of the baseline methods, in this scenario, NOTEARS-pool. We look for plausible explanations by investigating the results of the HSIC test in Table I. We conclude that the inferior performance of CICME is caused by the false identification of $X_{3}$ as unstable in most cases. If the causal mechanism of $X_{3}$ is not carried over to step 33, the causal edge $X_{2}\rightarrow X_{3}$ cannot be recovered because there is no variation in $X_{2}$ in each individual domain. This also partially explains the similar performance of both variants of CICME in this scenario.
@@ IV-E2 Execution Time
We compare the overall execution time of different methods in Fig. 4(a) based on (E1). Surprisingly, the execution time is not always positively correlated with the sample size. For instance, NOTEARS-pool and NOTEARS-ind take longer to run on 10 samples than on 100 samples. This can be attributed to the fact that the model sometimes requires more epochs to reach convergence when trained on fewer samples. Among all methods, NOTEARS-pool is most time efficient since it only involves the optimization of NOTEARS-MLP once. Interestingly, by freezing the weights of stable MLPs, CICME-f requires less execution time than NOTEARS-ind and, for small sample sizes, its execution time is only slightly higher than that of NOTEARS-pool, despite of the three-step approach. As expected, CICME-l takes the longest time due to the computational overhead created by the additional loss term. A closer look at the execution time of each step of CICME in Fig. 4(b) reveals that the execution time of HSIC test (step 22) is almost negligible, whereas applying NOTEARS-MLP takes relatively longer time. Freezing weights of stable MLPs greatly facilitates the estimation of the remaining causal mechanisms, indicated by the difference between step 33 in Fig. 4(b) and NOTEARS-ind in Fig. 4(a), further supporting the computational efficiency of CICME-f.
@@
Figure 4: Comparison of the execution time with different sample sizes. Step 33 (f) and Step 33 (l) in Fig 4(b) denotes the execution time of step 33 of CICME-f and CICME-l, respectively. The box plot was chosen over the violin plot in Fig. 4(b) for better color visualization.

### v4

@@ II-B Continuous Optimization-based Methods
$\begin{gathered}\min_{\theta}\quad\frac{1}{n}\sum_{j=1}^{d}\ell(\mathbf{x}_{j},\mathrm{MLP}(\mathbf{X};\theta_{j}))+\lambda\|A_{j}^{(1)}\|_{1}\\ \text{subject to }h(W(\theta)):=\mathrm{tr}(e^{W(\theta)\circ W(\theta)})-d=0~,\end{gathered}$ (7)
where $\ell$ is a loss function such as the least squares or the negative log-likelihood, $\ell_{1}$ regularization $\|A_{j}^{(1)}\|_{1}$ is used to enforce sparsity in the learned DAG, and the acyclicity constraint is enforced by $h(W(\theta))$ where $\circ$ denotes the Hadamard product and $e^{A}$ is the matrix exponential of $A$. Using the machinery of augmented Lagrangian method [24], the solution of the equality-constrained problem in (7) can be well approximated by solving an unconstrained $\ell_{1}$-penalized smooth minimization problem with the L-BFGS-B algorithm [25].
@@ III Methodology
III-A Step $1$: Causal Discovery on Pooled Data
@@ 4   Update $\theta$ to minimize the loss
III-B Step $2$: Detection of Stable Variables
@@ 8      Add node $X_{j}$ to $X_{\mathrm{stable}}$
III-C Step $3$: Causal Discovery on Individual Datasets
@@ IV Experiments
In this section, we generate synthetic data comprised of multiple domains based on linear Gaussian models to compare the performance of CICME and baseline methods. To assess CICME’s ability to detect common causal mechanisms, we consider $4$ scenarios inspired from a real manufacturing process where causal mechanisms may shift across domains.
@@ IV-A Baselines
  • Applying NOTEARS-MLP on the pooled dataset, referred to as NOTEARS-pool. It is equivalent to step $1$ of CICME.
  • Applying NOTEARS-MLP in each domain individually, referred to as NOTEARS-ind. It is equivalent to step $3$ of CICME with $X_{\mathrm{stable}}$ being an empty set.
@@ IV-E Results
We first examine the capability of CICME to identify common causal mechanisms across all domains based on the HSIC test results from step $2$. Note that in (E1) and (E2), $X_{4}$ exhibits changing causal mechanisms due to varying values of $H$ across domains, while in (E3) and (E4), $X_{2}$ is the expected unstable variable due to different distributions of $N_{2}$ across domains. Table I summarizes the total number of cases (out of $100$) in which each variable is identified as a stable variable. Additionally, the LSHD is averaged over those cases to indicate the accuracy of the estimated local structure. It can be seen that, when provided with $1000$ samples, CICME is able to reliably identify variables with changing causal mechanisms, as evidenced by the rare cases of misclassifications highlighted in bold. However, they become more frequent with a reduced sample size, since fewer samples provide less information about the population, leading to less statistical power of the HSIC test. Similarly, the accuracy of the estimated causal parents of stable variables also deteriorates when provided with fewer samples, as illustrated by the higher LSHD. This may adversely impact the performance of CICME by carrying over inaccurate common causal mechanisms to step $3$.
In addition, we also observe that stable variables are occasionally identified as unstable. While this leads to higher degrees of freedom in updating the weights of the model, and hence, more training effort in step $3$, it typically does not affect the overall performance. However, an exception is observed in (E4), where $X_{3}$ is identified as unstable significantly more often than stable even for $1000$ samples, despite the ground truth causal mechanism of $X_{3}$ being the same across all domains (i.e., $X_{3}=X_{1}+X_{2}+N_{3}$). We suspect that this is due to convergence issues in the constrained loss function. This can be hypothetically solved by finetuning the MLP based on the least squares loss before applying the HSIC test in step $2$.
@@ IV-E1 Causal Structure
Moreover, Fig. 3 compares the capability of all methods to recover causal structures when the causal mechanism of $X_{2}$ changes across domains. When $X_{2}$ follows a distribution with shifted means across domains, the performance of CICME is on par with either of the baseline methods, as shown in Fig. 3(a). Similarly, NOTEARS-pool again outperforms NOTEARS-ind for $10$ samples, for the same reason mentioned earlier. However, this trend changes as soon as $X_{2}$ is fixed in each domain. Fig. 3(b) illustrates that NOTEARS-pool significantly outperforms NOTEARS-ind regardless of the sample size. This expected behavior can be attributed to the variance of $X_{2}$ in the observed data: The variance of $X_{2}$ equals to zero in each individual dataset and becomes non-zero in the pooled dataset, making it possible to infer the causal relationship between $X_{2}$ and other variables. Due to the pooling operation in step $1$, CICME is also able to outperform NOTEARS-ind regardless of the sample size, and it even achieves the best performance among all when the sample size is $10$. However, for a sample size of $1000$, CICME fails to match the performance of the best of the baseline methods, in this scenario, NOTEARS-pool. We look for plausible explanations by investigating the results of the HSIC test in Table I. We conclude that the inferior performance of CICME is caused by the false identification of $X_{3}$ as unstable in most cases. If the causal mechanism of $X_{3}$ is not carried over to step $3$, the causal edge $X_{2}\rightarrow X_{3}$ cannot be recovered because there is no variation in $X_{2}$ in each individual domain. This also partially explains the similar performance of both variants of CICME in this scenario.
@@ IV-E2 Execution Time
We compare the overall execution time of different methods in Fig. 4(a) based on (E1). Surprisingly, the execution time is not always positively correlated with the sample size. For instance, NOTEARS-pool and NOTEARS-ind take longer to run on 10 samples than on 100 samples. This can be attributed to the fact that the model sometimes requires more epochs to reach convergence when trained on fewer samples. Among all methods, NOTEARS-pool is most time efficient since it only involves the optimization of NOTEARS-MLP once. Interestingly, by freezing the weights of stable MLPs, CICME-f requires less execution time than NOTEARS-ind and, for small sample sizes, its execution time is only slightly higher than that of NOTEARS-pool, despite of the three-step approach. As expected, CICME-l takes the longest time due to the computational overhead created by the additional loss term. A closer look at the execution time of each step of CICME in Fig. 4(b) reveals that the execution time of HSIC test (step $2$) is almost negligible, whereas applying NOTEARS-MLP takes relatively longer time. Freezing weights of stable MLPs greatly facilitates the estimation of the remaining causal mechanisms, indicated by the difference between step $3$ in Fig. 4(b) and NOTEARS-ind in Fig. 4(a), further supporting the computational efficiency of CICME-f.
@@
Figure 4: Comparison of the execution time with different sample sizes. Step 33 (f) and Step 33 (l) in Fig 4(b) denotes the execution time of step $3$ of CICME-f and CICME-l, respectively. The box plot was chosen over the violin plot in Fig. 4(b) for better color visualization.